# General Order Virtual Element Methods for Neumann Boundary Optimal Control Problems in Saddle Point Formulation

### v1

@@ 3. VEM Numerical Discretization
(20) ${m}^{E}_{h}\left(w_{h},q_{h}\right)\vcentcolon=\left(\Pi^{0,E}_{k}w_{h},\Pi^{0,E}_{k}q_{h}\right)_{E}\,,$
for each $w_{h},q_{h}\in Y_{h}(E)$. The bilinear form $S^{E}:Y_{h}(E)\times Y_{h}(E)\to\mathbb{R}$, called stabilization, is any symmetric positive definite bilinear form to be chosen to verify that $\exists s_{\ast},s^{\ast}>0$ independent of $h_{E}$ such that
(21) $s_{\ast}\left(\nabla w_{h},\nabla w_{h}\right)_{E}\leq S^{E}\left(w_{h},w_{h}\right)\leq s^{\ast}\left(\nabla w_{h},\nabla w_{h}\right)_{E},$
for each $w_{h}\in Y_{h}(E)$ verifying $\Pi^{\nabla,E}_{k}w_{h}=0$.
Remark 2.
A common choice for $S^{E}$ is the dofi-dofi stabilizing bilinear form $S^{E}\colon Y_{h}(E)\times Y_{h}(E)\to\mathbb{R}$, defined such that, denoting by $\chi^{E}(v_{h})$ the vector of degrees of freedom of $v_{h}$ on $E$ and by $\sigma$ an arbitrary-chosen positive constant,
(22) $S^{E}\left(u_{h},v_{h}\right)=\sigma\;\chi^{E}(u_{h})\cdot\chi^{E}(v_{h})\quad\forall u_{h},v_{h}\in Y_{h}(E)\,.$
We report the following classical result [4].
@@
There exists $\eta_{\ast}\,,\eta^{\ast}\,,\mu^{\ast}>0$, independent of $h_{E}$, such that for each $w_{h}\in Y_{h}(E)$
(23) $\eta_{\ast}\left\lvert\left\lvert\left\lvert w_{h}\right\rvert\right\rvert\right\rvert_{Y(E)}^{2}\leq{a}^{E}_{h}\left(w_{h},w_{h}\right)\leq\eta^{\ast}\left\lvert\left\lvert\left\lvert w_{h}\right\rvert\right\rvert\right\rvert_{Y(E)}^{2}$
and
(24) ${m}^{E}_{h}\left(w_{h},w_{h}\right)\leq\mu^{\ast}\left\lVert{w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2}\,.$
We define the global bilinear forms such that for each $w_{h},q_{h}\in Y_{h}$,
${a}_{h}\left(w_{h},q_{h}\right)\vcentcolon=\sum_{E\in\mathcal{M}_{h}}{a}^{E}_{h}\left(w_{h},q_{h}\right)\text{ and }{m}_{h}\left(w_{h},q_{h}\right)\vcentcolon=\sum_{E\in\mathcal{M}_{h}\cap\Omega_{\text{obs}}}{m}^{E}_{h}\left(w_{h},p_{h}\right)\,.$
Let us now write the VEM approximation of Problem 3. Let us consider $\boldsymbol{x}_{h}\vcentcolon=(y_{h},u_{h})\in X_{h}$, where $X_{h}\vcentcolon=Y_{h}\times U_{h}\subset X$ equipped with the norm $\left\lVert{\cdot}\right\rVert_{X}$, i.e. $\left\lVert{\boldsymbol{x}_{h}}\right\rVert_{X}^{2}=\left\lvert\left\lvert\left\lvert y_{h}\right\rvert\right\rvert\right\rvert_{Y}^{2}+\left\lVert{u_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}^{2}$ , and the bilinear forms ${\mathcal{A}}_{h}:X_{h}\times X_{h}\to\mathbb{R}$ and ${\mathcal{B}}_{h}:Y_{h}\times X_{h}\to\mathbb{R}$ given by
@@
${\mathcal{B}}_{h}\left(q_{h},\boldsymbol{x}_{h}\right)={a}_{h}\left(y_{h},q_{h}\right)-\left(u_{h},q_{h}\right)_{\Gamma_{C}}\;\forall\boldsymbol{x}_{h}\in X_{h},\,\forall q_{h}\in Y_{h}\,.$
@@
$_{h}{}{X_{h}^{*},X_{h}}=\sum_{E\in\mathcal{M}_{h}\cap\Omega_{\text{obs}}}\left(y_{\mathrm{d}},\Pi^{0,E}_{k}w_{h}\right)_{E}+\left(0,v_{h}\right)_{\Gamma_{C}}\,.$
@@
    (26) ${\mathcal{A}}_{h}\left(\boldsymbol{\varsigma}_{h},\boldsymbol{\varsigma}_{h}\right)\geq\alpha_{\ast}\,\left\lVert{\boldsymbol{\varsigma}_{h}}\right\rVert_{X}^{2}\quad\forall\boldsymbol{\varsigma}_{h}\in X_{h,0}.$
@@
    (27) $\inf_{p_{h}\in Y_{h}}\sup_{\boldsymbol{\varsigma}_{h}\in X_{h}}\frac{{\mathcal{B}}_{h}\left(p_{h},\boldsymbol{\varsigma}_{h}\right)}{\left\lvert\left\lvert\left\lvert p_{h}\right\rvert\right\rvert\right\rvert_{Y}\left\lVert{\boldsymbol{\varsigma}_{h}}\right\rVert_{X}}\geq\beta_{\ast}\,.$
@@
(28) $J_{h}(y_{h},u_{h})=\frac{1}{2}\sum_{E\in\mathcal{M}_{h}\cap\Omega_{\text{obs}}}\left\lVert{\Pi^{0,E}_{k}y_{h}-y_{\mathrm{d}}}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2}+\frac{\alpha}{2}\left\lVert{u_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}^{2},$
@@
$|{\mathcal{A}}_{h}\left(\boldsymbol{x}_{h},\boldsymbol{\varsigma}_{h}\right)|$ $\leq\sum_{E\in\mathcal{M}_{h}\cap\Omega_{\text{obs}}}\left\lVert{\Pi^{0,E}_{k}y_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}\left\lVert{\Pi^{0,E}_{k}q_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}+\alpha\left\lVert{u_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}\left\lVert{v_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}$
@@
$\leq C\left(1+\alpha\right)\left\lVert{\boldsymbol{x}_{h}}\right\rVert_{X}\left\lVert{\boldsymbol{\varsigma}_{h}}\right\rVert_{X}\,,$
@@
Using Lemma 1, summing over all the elements of $\mathcal{M}_{h}$ and calling $\eta_{\ast}$ the global coercivity constant with abuse of notation, applying the trace inequality $\left\lVert{w_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}\leq C_{T}\left\lvert\left\lvert\left\lvert w_{h}\right\rvert\right\rvert\right\rvert_{Y}$ we obtain
(29) $\eta_{\ast}\left\lvert\left\lvert\left\lvert w_{h}\right\rvert\right\rvert\right\rvert_{Y}^{2}\leq{a}_{h}\left(w_{h},w_{h}\right)=\left(v_{h},w_{h}\right)_{\Gamma_{C}}$ $\leq\left\lVert{v_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}\left\lVert{w_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}$
@@
(30) $\left\lvert\left\lvert\left\lvert w_{h}\right\rvert\right\rvert\right\rvert_{Y}\leq\frac{C_{T}}{\eta_{\ast}}\left\lVert{v_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}.$
@@
${\mathcal{A}}_{h}\left(\boldsymbol{\varsigma}_{h},\boldsymbol{\varsigma}_{h}\right)$ $=\sum_{E\in\mathcal{M}_{h}\cap\Omega_{\text{obs}}}\left\lVert{\Pi^{0,E}_{k}w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2}+\alpha\left\lVert{v_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}^{2}$
$=\sum_{E\in\mathcal{M}_{h}\cap\Omega_{\text{obs}}}\left\lVert{\Pi^{0,E}_{k}w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2}+\frac{\alpha}{2}\left\lVert{v_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}^{2}+\frac{\alpha}{2}\left\lVert{v_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}^{2}$
@@
4. Error Estimates
In order to derive optimal a priori error estimates, we have to provide the polynomial approximation properties and the interpolation error estimates, presented in Lemmas 2, 3 and 4, respectively. The results about polynomial approximation can be led back to classical results ([14]), while the VEM interpolation error estimate is proved in [17] and the interpolation on $U_{h}$ derives from standard results of one-dimensional FEM interpolation.
@@ 4. Error Estimates
(31) $\left\lvert{\phi-\Pi^{0,E}_{k}\phi}\right\rvert_{m,E}\leq Ch_{E}^{s-m}\left\lvert{\phi}\right\rvert_{s,E},\;\mbox{ with }m\leq s\leq k+1\,,$
@@
(32) $\left\lvert{\phi-\Pi^{\nabla,E}_{k}\phi}\right\rvert_{m,E}\leq Ch_{E}^{s-m}\left\lvert{\phi}\right\rvert_{s,E},\;\mbox{ with }m\leq s\leq k+1\,.$
@@
Under assumptions A.1 and A.2, let $w\in\mathrm{H}^{s+1}\!(\Omega)$, $0\leq s\leq k$, there exists $C>0$ such that $\forall\,h$, $\exists w_{I}\in Y_{h}$ satisfying
(33) $\left\lVert{w-w_{I}}\right\rVert_{\mathrm{L}^{2}\!(\Omega)}+h\left\lVert{\nabla(w-w_{I})}\right\rVert_{\mathrm{L}^{2}\!(\Omega)}\leq Ch^{s+1}\left\lvert{w}\right\rvert_{s+1,\Omega},$
where $C$ is a positive constant, independent of $h$.
@@
Under assumptions A.1 and A.2, let $v\in\mathrm{H}^{r}\!(\Gamma_{C})$, $10\leq r\leq k+$ 1. Then there exists a function $v_{I}\in U_{h}$ and a constant $C>0$ independent of $h$ such that
(34) $\left\lVert{v-v_{I}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}\leq Ch^{r}\left\lvert{v}\right\rvert_{r,\Gamma_{C}}\,.$
@@
(35) ${\mathcal{B}}_{h}\left(q_{h},\left(y_{{\mathcal{B}}_{h}}-y_{h},u_{I}-u_{h}\right)\right)=0\quad\forall q_{h}\in Y_{h}\,.$
Moreover, if $y\in\mathrm{H}^{s+1}\!(\Omega)$, $u\in\mathrm{H}^{r}\!(\Gamma_{C})$, $f\in\mathrm{H}^{t}\!(\Omega)$ with $0\leq s\leq k$, $10\leq r\leq k+$ 1 and $10\leq t\leq k+$ 1, then there exists $C>0$ independent of $h$ such that
(36) $\left\lvert\left\lvert\left\lvert y-y_{{\mathcal{B}}_{h}}\right\rvert\right\rvert\right\rvert_{Y}\leq C\left(h^{s}\left\lvert{y}\right\rvert_{s+1,\Omega}+h^{t+1}\left\lvert{f}\right\rvert_{t,\Omega}+h^{r}\left\lvert{u}\right\rvert_{r,\Gamma_{C}}\right)\,.$
@@
(37) ${a}_{h}\left(y_{{\mathcal{B}}_{h}},q_{h}\right)=\left(u_{I},q_{h}\right)_{\Gamma_{C}}+\sum_{E\in\mathcal{M}_{h}}\left(\Pi^{0,E}_{k}f,q_{h}\right)\quad\forall q_{h}\in Y_{h}\,.$
Notice that, by definition of ${\mathcal{B}}_{h}$ (3), we have
${\mathcal{B}}_{h}\left(q_{h},(y_{{\mathcal{B}}_{h}},u_{I})\right)=\sum_{E\in\mathcal{M}_{h}}\left(\Pi^{0,E}_{k}f,q_{h}\right)={\mathcal{B}}_{h}\left(q_{h},(y_{h},u_{h})\right)\,,$
@@
$\left\lvert\left\lvert\left\lvert y-y_{{\mathcal{B}}_{h}}\right\rvert\right\rvert\right\rvert_{Y}\leq\left\lvert\left\lvert\left\lvert y-y_{I}\right\rvert\right\rvert\right\rvert_{Y}+\left\lvert\left\lvert\left\lvert y_{I}-y_{{\mathcal{B}}_{h}}\right\rvert\right\rvert\right\rvert_{Y}\leq Ch^{s}\left\lvert{y}\right\rvert_{s+1,\Omega}+\left\lvert\left\lvert\left\lvert y_{I}-y_{{\mathcal{B}}_{h}}\right\rvert\right\rvert\right\rvert_{Y}\,,$
where $C>0$ is independent of $h$. Regarding the second term above, let $e_{h}=y_{I}-y_{{\mathcal{B}}_{h}}$. Exploiting (23), (37) and the second equation of Problem 3 we get
(38) $\begin{split}\left\lvert\left\lvert\left\lvert y_{I}-y_{{\mathcal{B}}_{h}}\right\rvert\right\rvert\right\rvert_{Y}^{2}&\leq C{a}_{h}\left(y_{I}-y_{{\mathcal{B}}_{h}},e_{h}\right)\\ &=C\Big{(}{a}_{h}\left(y_{I},e_{h}\right)-{a}\left(y,e_{h}\right)\\ &\qquad\qquad+\sum_{E\in\mathcal{M}_{h}}\left(f-\Pi^{0,E}_{k}f,e_{h}\right)_{E}+\left(u-u_{I},e_{h}\right)_{\Gamma_{C}}\Big{)}\,.\end{split}$
The terms above are estimated separately, as follows. By the continuity of ${a}$ and ${a}_{h}$, a triangle inequality and estimates (33) and (32), we get
(39) $\begin{split}{a}_{h}\left(y_{I},e_{h}\right)-{a}\left(y,e_{h}\right)&=\sum_{E\in\mathcal{M}_{h}}{a}^{E}_{h}\left(y_{I}-\Pi^{0,E}_{k}y,e_{h}\right)+{a}^{E}_{h}\left(\Pi^{0,E}_{k}y,e_{h}\right)\\ &\qquad-{a}^{E}\left(\Pi^{0,E}_{k}y,e_{h}\right)+{a}^{E}\left(\Pi^{0,E}_{k}y-y,e_{h}\right)\\ &\leq\sum_{E\in\mathcal{M}_{h}}{a}^{E}_{h}\left(\Pi^{0,E}_{k}y,e_{h}\right)-{a}^{E}\left(\Pi^{0,E}_{k}y,e_{h}\right)\\ &\qquad+C\Big{(}\sum_{E\in\mathcal{M}_{h}}\left\lvert\left\lvert\left\lvert y-\Pi^{0,E}_{k}y\right\rvert\right\rvert\right\rvert_{Y(E)}\\ &\qquad\qquad\qquad+\left\lvert\left\lvert\left\lvert y-y_{I}\right\rvert\right\rvert\right\rvert_{Y(E)}\Big{)}\left\lvert\left\lvert\left\lvert e_{h}\right\rvert\right\rvert\right\rvert_{Y}\\ &\leq\sum_{E\in\mathcal{M}_{h}}{a}^{E}_{h}\left(\Pi^{0,E}_{k}y,e_{h}\right)-{a}^{E}\left(\Pi^{0,E}_{k}y,e_{h}\right)\\ &\qquad\qquad\qquad\qquad\qquad+Ch^{s}\left\lvert{y}\right\rvert_{s+1,\Omega}\left\lvert\left\lvert\left\lvert e_{h}\right\rvert\right\rvert\right\rvert_{Y}\,.\end{split}$
Moreover, the following estimate can be proved by proceeding as in [4, Lemma 5.5]:
$\sum_{E\in\mathcal{M}_{h}}{a}^{E}_{h}\left(\Pi^{0,E}_{k}y,e_{h}\right)-{a}^{E}\left(\Pi^{0,E}_{k}y,e_{h}\right)\leq Ch^{s}\left\lvert{y}\right\rvert_{s+1,\Omega}\left\lvert\left\lvert\left\lvert e_{h}\right\rvert\right\rvert\right\rvert_{Y}\,.$

### v2

@@ 3 Numerical Discretization
${m}^{E}_{h}\left(w_{h},q_{h}\right)\vcentcolon=\left(\Pi^{0,E}_{k}w_{h},\Pi^{0,E}_{k}q_{h}\right)_{E}.$ (24)
We define the global bilinear forms such that for each $w_{h},q_{h}\in\mathrm{H}^{1}\!(\Omega)$,
${a}_{h}\left(w_{h},q_{h}\right)\vcentcolon=\sum_{E\in\mathcal{M}_{h}}{a}^{E}_{h}\left(w_{h},q_{h}\right)\text{ and }{m}_{h}\left(w_{h},q_{h}\right)\vcentcolon=\sum_{E\in\mathcal{M}_{h,\text{obs}}}{m}^{E}_{h}\left(w_{h},p_{h}\right)\,.$
To assess the coercivity and the continuity of the bilinear forms in the SFVEM approximation, we recall the following assumption from [5]
  1. A.3
    For any $E\in\mathcal{M}_{h}$, let $\ell_{E}$ be the smallest integer such that any polynomial $\pi_{k+\ell_{E}}\in\mathcal{P}_{k+\ell_{E}}(E)$ can be identified by a set of degrees of freedom which contains $kN_{E}-1$ distinct moments $\frac{1}{|\partial E|}\,(\pi_{k+\ell_{E}},\xi_{i})_{\partial E},$ for a scaled polynomial basis of $\mathbb{P}^{0}_{k-1}\!\left(\partial E\right)\coloneqq\left\{\xi\;:\;\xi|_{e}\in\mathbb{P}_{k-1}(e),\ \forall e\subset\partial E,\ \int_{\partial E}\xi=0\right\}.$
We report the following result, easily derivable from [5], concerning the coercivity and the continuity of the bilinear form $a_{h}$, and from [3], concerning the continuity of $m_{h}$.
@@
Under the assumptions A.1, A.2, for each $w_{h}\in\mathrm{H}^{1}\!(\Omega)$
${m}_{h}\left(w_{h},w_{h}\right)\leq\left\lVert{w_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Omega)}^{2}\,$, (25)
and there exists $\eta^{\ast}>0$ such that
${a}_{h}\left(w_{h},w_{h}\right)\leq\eta^{\ast}\left\lvert\left\lvert\left\lvert w_{h}\right\rvert\right\rvert\right\rvert_{Y}^{2}\,.$ (26)
Assuming also A.3, for each $w_{h}\in Y_{h}$ there exists $\eta_{\ast}>0$, such that, for $h$ sufficiently small,
${a}_{h}\left(w_{h},w_{h}\right)\geq\eta_{\ast}\left\lvert\left\lvert\left\lvert w_{h}\right\rvert\right\rvert\right\rvert_{Y}^{2}\,.$ (27)
Remark 2 (Standard VEM approximation).
We here highlight the main difference between the proposed approximation and the one involving a standard VEM scheme (see [3]). For each $E\in\mathcal{M}_{h}$, let ${\tilde{a}}^{E}_{h}:Y_{h}(E)\times Y_{h}(E)\to\mathbb{R}$ be the local bilinear form defined as
${\tilde{a}}^{E}_{h}\left(w_{h},q_{h}\right)\vcentcolon=\left(\mathcal{K}\Pi^{0,E}_{k-1}\nabla w_{h},\Pi^{0,E}_{k-1}\nabla q_{h}\right)_{E}\\ +\left\lVert{\mathcal{K}}\right\rVert_{\infty,E}S^{E}\left(w_{h}-\Pi^{\nabla,E}_{k}w_{h},q_{h}-\Pi^{\nabla,E}_{k}q_{h}\right)\\ +\left(\beta\Pi^{0,E}_{k-1}\nabla w_{h},\Pi^{0,E}_{k}q_{h}\right)_{E}+\left(\gamma\Pi^{0,E}_{k}w_{h},\Pi^{0,E}_{k}q_{h}\right)_{E}.$ (28)
The bilinear form $S^{E}:Y_{h}(E)\times Y_{h}(E)\to\mathbb{R}$, called stabilization, is any symmetric positive definite bilinear form to be chosen to verify that $\exists s_{\ast},s^{\ast}>0$ independent of $h_{E}$ such that
$s_{\ast}\left(\nabla w_{h},\nabla w_{h}\right)_{E}\leq S^{E}\left(w_{h},w_{h}\right)\leq s^{\ast}\left(\nabla w_{h},\nabla w_{h}\right)_{E},$ (29)
for each $w_{h}\in Y_{h}(E)$ verifying $\Pi^{\nabla,E}_{k}w_{h}=0$. A common choice for $S^{E}$ is the dofi-dofi stabilizing bilinear form $S^{E}\colon Y_{h}(E)\times Y_{h}(E)\to\mathbb{R}$, defined such that, denoting by $\chi^{E}(v_{h})$ the vector of degrees of freedom of $v_{h}$ on $E$ and by $\sigma$ an arbitrary-chosen positive constant,
$S^{E}\left(w_{h},q_{h}\right)=\sigma\;\chi^{E}(w_{h})\cdot\chi^{E}(q_{h})\quad\forall w_{h},q_{h}\in Y_{h}(E)\,.$ (30)
We recall that other choices of $S^{E}$ can be made, allowing a relaxation of the mesh assumption A.2 [13]. For each $w_{h},q_{h}\in Y_{h}$, we now define the global VEM form
$\tilde{a}_{h}(w_{h},q_{h})\vcentcolon=\sum_{E\in\mathcal{M}_{h}}{\tilde{a}}^{E}_{h}\left(w_{h},q_{h}\right)\,$,
that satisfies under the assumptions A.1 and A.2, there exists $\tilde{\eta}_{\ast}\,,\tilde{\eta}^{\ast}>0$, independent of $h_{E}$, such that for each $w_{h}\in Y_{h}$
$\tilde{\eta}_{\ast}\left\lvert\left\lvert\left\lvert w_{h}\right\rvert\right\rvert\right\rvert_{Y}^{2}\leq\tilde{a}_{h}(w_{h},w_{h})\leq\tilde{\eta}^{\ast}\left\lvert\left\lvert\left\lvert w_{h}\right\rvert\right\rvert\right\rvert_{Y}^{2}.$ (31)
Let us now write the proposed approximation of Problem 3. Let us consider $\boldsymbol{x}_{h}\vcentcolon=(y_{h},u_{h})\in X_{h}$, where $X_{h}\vcentcolon=Y_{h}\times U_{h}\subset X$ equipped with the norm $\left\lVert{\cdot}\right\rVert_{X}$, i.e. $\left\lVert{\boldsymbol{x}_{h}}\right\rVert_{X}^{2}=\left\lvert\left\lvert\left\lvert y_{h}\right\rvert\right\rvert\right\rvert_{Y}^{2}+\left\lVert{u_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}^{2}$ , and the bilinear forms ${\mathcal{A}}_{h}:X_{h}\times X_{h}\to\mathbb{R}$ and ${\mathcal{B}}_{h}:Y_{h}\times X_{h}\to\mathbb{R}$ given by
@@
${\mathcal{B}}_{h}\left(q_{h},\boldsymbol{x}_{h}\right)={a}_{h}\left(y_{h},q_{h}\right)-\left(u_{h},q_{h}\right)_{\Gamma_{C}}\;\forall\boldsymbol{x}_{h}\in X_{h},\,\forall q_{h}\in Y_{h}\,.$ (32)
@@
[…]
@@
    ${\mathcal{A}}_{h}\left(\boldsymbol{\varsigma}_{h},\boldsymbol{\varsigma}_{h}\right)\geq\alpha_{\ast}\,\left\lVert{\boldsymbol{\varsigma}_{h}}\right\rVert_{X}^{2}\quad\forall\boldsymbol{\varsigma}_{h}\in X_{h,0}.$ (34)
@@
    $\inf_{p_{h}\in Y_{h}}\sup_{\boldsymbol{\varsigma}_{h}\in X_{h}}\frac{{\mathcal{B}}_{h}\left(p_{h},\boldsymbol{\varsigma}_{h}\right)}{\left\lvert\left\lvert\left\lvert p_{h}\right\rvert\right\rvert\right\rvert_{Y}\left\lVert{\boldsymbol{\varsigma}_{h}}\right\rVert_{X}}\geq\beta_{\ast}\,.$ (35)
@@
$J_{h}(y_{h},u_{h})=\frac{1}{2}\sum_{E\in\mathcal{M}_{h,\text{obs}}}\left\lVert{\Pi^{0,E}_{k}y_{h}-y_{\mathrm{d}}}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2}+\frac{\alpha}{2}\left\lVert{u_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}^{2},$ (36)
@@
$|{\mathcal{A}}_{h}\left(\boldsymbol{x}_{h},\boldsymbol{\varsigma}_{h}\right)|$ $\leq\sum_{E\in\mathcal{M}_{h,\text{obs}}}\left\lVert{\Pi^{0,E}_{k}y_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}\left\lVert{\Pi^{0,E}_{k}q_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}+\alpha\left\lVert{u_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}\left\lVert{v_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}$
@@
$\leq{C^{*}}\left(1+\alpha\right)\left\lVert{\boldsymbol{x}_{h}}\right\rVert_{X}\left\lVert{\boldsymbol{\varsigma}_{h}}\right\rVert_{X}\,$,
@@
Using Lemma 1, applying the discrete trace inequality (see e.g. [17]) and Remark 1, $\left\lVert{w_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}\leq C_{T}\left\lvert\left\lvert\left\lvert w_{h}\right\rvert\right\rvert\right\rvert_{Y}$ we obtain
$\eta_{\ast}\left\lvert\left\lvert\left\lvert w_{h}\right\rvert\right\rvert\right\rvert_{Y}^{2}\leq{a}_{h}\left(w_{h},w_{h}\right)=\left(v_{h},w_{h}\right)_{\Gamma_{C}}$ $\leq\left\lVert{v_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}\left\lVert{w_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}$ (37)
@@
$\left\lvert\left\lvert\left\lvert w_{h}\right\rvert\right\rvert\right\rvert_{Y}\leq\frac{C_{T}}{\eta_{\ast}}\left\lVert{v_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}.$ (38)
@@
${\mathcal{A}}_{h}\left(\boldsymbol{\varsigma}_{h},\boldsymbol{\varsigma}_{h}\right)$ $=\sum_{E\in\mathcal{M}_{h,\text{obs}}}\left\lVert{\Pi^{0,E}_{k}w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2}+\alpha\left\lVert{v_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}^{2}$
$=\sum_{E\in\mathcal{M}_{h,\text{obs}}}\left\lVert{\Pi^{0,E}_{k}w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2}+\frac{\alpha}{2}\left\lVert{v_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}^{2}+\frac{\alpha}{2}\left\lVert{v_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}^{2}$
@@
4 Error Estimates
In order to derive optimal a priori error estimates, we have to provide the polynomial approximation properties and the interpolation error estimates, presented in Lemmas 2, 3 and 4, respectively. The results about polynomial approximation can be led back to classical results, while the VEM interpolation error estimate is proved in [24] and the interpolation on $U_{h}$ derives from standard results of one-dimensional FEM interpolation.
@@ 4 Error Estimates
$\left\lvert{\phi-\Pi^{0,E}_{k}\phi}\right\rvert_{m,E}\leq Ch_{E}^{s-m}\left\lvert{\phi}\right\rvert_{s,E},\;\mbox{ with }m\leq s\leq k+1\,$, (39)
@@
$\left\lvert{\phi-\Pi^{\nabla,E}_{k}\phi}\right\rvert_{m,E}\leq Ch_{E}^{s-m}\left\lvert{\phi}\right\rvert_{s,E},\;\mbox{ with }m\leq s\leq k+1\,.$ (40)
@@
Under assumptions A.1 and A.2, let $w\in\mathrm{H}^{s+1}\!(\Omega)$, $0\leq s\leq k$, there exists $C>0$, independent of $h$, such that $\forall\,h$, $\exists w_{I}\in Y_{h}$ satisfying
$\left\lVert{w-w_{I}}\right\rVert_{\mathrm{L}^{2}\!(\Omega)}+h\left\lVert{\nabla(w-w_{I})}\right\rVert_{\mathrm{L}^{2}\!(\Omega)}\leq Ch^{s+1}\left\lvert{w}\right\rvert_{s+1,\Omega}.$ (41)
Thus,
$\left\lvert\left\lvert\left\lvert w-w_{I}\right\rvert\right\rvert\right\rvert_{Y}\leq Ch^{s}\left\lvert{w}\right\rvert_{s+1,\Omega}.$ (42)
@@
Under assumptions A.1 and A.2, let
$v\in\mathrm{H}^{r}\!(\Gamma_{C})$, $0\leq r\leq k+1$. Then there exists a function $v_{I}\in U_{h}$ and a constant $C>0$ independent of $h$ such that
$\left\lVert{v-v_{I}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}\leq Ch^{r}\left\lvert{v}\right\rvert_{r,\Gamma_{C}}\,.$ (43)
@@
${\mathcal{B}}_{h}\left(q_{h},\left(y_{{\mathcal{B}}_{h}}-y_{h},u_{I}-u_{h}\right)\right)=0\quad\forall q_{h}\in Y_{h}\,.$ (44)
Moreover, if $y\in\mathrm{H}^{s+1}\!(\Omega)$, $u\in\mathrm{H}^{r}\!(\Gamma_{C})$, $f\in\mathrm{H}^{t}\!(\Omega)$ with $0\leq s\leq k$, $0\leq r\leq k+1$ and $0\leq t\leq k+1$, then there exists $C>0$ independent of $h$ such that
$\left\lvert\left\lvert\left\lvert y-y_{{\mathcal{B}}_{h}}\right\rvert\right\rvert\right\rvert_{Y}\leq C\left(h^{s}\left\lvert{y}\right\rvert_{s+1,\Omega}+h^{t+1}\left\lvert{f}\right\rvert_{t,\Omega}+h^{r}\left\lvert{u}\right\rvert_{r,\Gamma_{C}}\right)\,.$ (45)
@@
${a}_{h}\left(y_{{\mathcal{B}}_{h}},q_{h}\right)=\left(u_{I},q_{h}\right)_{\Gamma_{C}}+\sum_{E\in\mathcal{M}_{h}}\left(f,\Pi^{0,E}_{k}q_{h}\right)\quad\forall q_{h}\in Y_{h}\,.$ (46)
Notice that, by definition of ${\mathcal{B}}_{h}$ (32) and Problem 4, we have
${\mathcal{B}}_{h}\left(q_{h},(y_{{\mathcal{B}}_{h}},u_{I})\right)=\sum_{E\in\mathcal{M}_{h}}\left(f,\Pi^{0,E}_{k}q_{h}\right)={\mathcal{B}}_{h}\left(q_{h},(y_{h},u_{h})\right)\,$,
@@
$\left\lvert\left\lvert\left\lvert y-y_{{\mathcal{B}}_{h}}\right\rvert\right\rvert\right\rvert_{Y}\leq\left\lvert\left\lvert\left\lvert y-y_{I}\right\rvert\right\rvert\right\rvert_{Y}+\left\lvert\left\lvert\left\lvert y_{I}-y_{{\mathcal{B}}_{h}}\right\rvert\right\rvert\right\rvert_{Y}\leq Ch^{s}\left\lvert{y}\right\rvert_{s+1,\Omega}+\left\lvert\left\lvert\left\lvert y_{I}-y_{{\mathcal{B}}_{h}}\right\rvert\right\rvert\right\rvert_{Y}\,.$
Regarding the second term above, let $e_{h}=y_{I}-y_{{\mathcal{B}}_{h}}$. Exploiting (27), (46) and the second equation of Problem 3 we get
$\begin{split}\left\lvert\left\lvert\left\lvert y_{I}-y_{{\mathcal{B}}_{h}}\right\rvert\right\rvert\right\rvert_{Y}^{2}&\leq C{a}_{h}\left(y_{I}-y_{{\mathcal{B}}_{h}},e_{h}\right)\\ &=C\Big({a}_{h}\left(y_{I},e_{h}\right)-{a}\left(y,e_{h}\right)\\ &\qquad\qquad+\sum_{E\in\mathcal{M}_{h}}\left(f-\Pi^{0,E}_{k}f,e_{h}\right)_{E}+\left(u-u_{I},e_{h}\right)_{\Gamma_{C}}\Big)\,.\end{split}$ (47)
The terms above are estimated separately, as follows. Adding and subtracting $\sum_{E\in\mathcal{M}_{h}}{a}^{E}_{h}\left(y,e_{h}\right)$, applying the continuity of ${a}_{h}$ (26) and the estimate (42), we get
[…]
where the last estimate can be derived as in [5, Theorem 5].